An object is thrown with an initial velocity of 14 ft per second. Its height is given by h= -14t^2 + 56 at time t seconds. After how many seconds does it hit the ground?
Answer by ikdeep(189)

we have...
speed = 14 ft/second
distance or h = -14t^2 + 56
also we know that ...Distance = speed * time
putting the above values in the formula..we get
-14t^2 + 56 = 14 * t
taking common 14 from LHS ,,we get
14 ( -t^2 + 4) = 14 * t
divide both sides by 14 and we get ...
-t^2 + 4 = t
or 0 = t^2 + t - 4
solve the value of t by discriminent and you get the value of t
and put the value of t in formula ..distance = speed * time
and you get distance also.,.
Still if you unable to solve further contact me again.
Hope this will help you
Please feel free to revert back for any further queries.